OpenStudy (anonymous):

Boyle's Law says that the volume of a gas varies inversely with the pressure. When the volume of a certain gas is 10L, the pressure is 195 kPa (kilopascals). What is the pressure when the volume is 13L ?

OpenStudy (doc.brown):

The pressure and volume are proportional.\[P\alpha\frac{1}{V}\]That means if the pressure goes up, the volume goes down. They do this at the same rate. Double the pressure, half the volume.\[\color{red}{2}P\alpha\frac{1}{\color{red}{2}V}\]I like to think of P and V together making some constant.\[PV=k\]Double the pressure, half the volume.\[\color{red}{2}P\frac{V}{\color{red}{2}}=k\]What you do to one you have to do the opposite to the other to keep them constant.

OpenStudy (doc.brown):

Increase the volume from 10L to 13L, you obviously have to decrease the pressure by the same proportion, from 195kPa to what?